OpenStudy (itsmichelle29):

If sin θ = 1/4 and tan θ > 0, what is the value of cos θ?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

draw a triangle, like this 

OpenStudy (misty1212):

there is a picture of an angle where the sine is \(\frac{1}{4}\) since the "opposite over adjacent" is \(\frac{1}{4}\)

OpenStudy (misty1212):

i meant "opposite over hypotenuse"

OpenStudy (misty1212):

what you need is the adjacent side, which you find via pythagoras let me know when you get it

OpenStudy (itsmichelle29):

7

OpenStudy (misty1212):

oh no not 7

OpenStudy (misty1212):

use pythagoras \[a^2+b^2=c^2\\ x^2+1^2=4^2\\ x^2+1=16\\ x^2=16-1=15\] so \[x=\sqrt{15}\]

OpenStudy (itsmichelle29):

okay is the answe root15 /4

OpenStudy (itsmichelle29):

@misty1212

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

oh no!

OpenStudy (misty1212):

you want the tangent right? not the cosine

OpenStudy (itsmichelle29):

can u help me with like two more plz

OpenStudy (itsmichelle29):

the cos

OpenStudy (misty1212):

oh ok then yes

OpenStudy (misty1212):

and i will be happy to help with two more

OpenStudy (itsmichelle29):

thanks

OpenStudy (itsmichelle29):

What are the period and phase shift for f(x) = −4 tan(x − π)?

OpenStudy (itsmichelle29):

i think its pi and pi

OpenStudy (misty1212):

period of tangent is \(\pi\) so yeah

OpenStudy (misty1212):

phase shift is \(\pi\) as well you are right

OpenStudy (itsmichelle29):

thanks and

OpenStudy (itsmichelle29):

Use the graph below to answer the question that follows: cosine graph with points at 0, negative 1 and pi over 2, 3 and pi, negative 1 What are the amplitude, period, and midline of the function? Amplitude: 8; period: π; midline: y = 1 Amplitude: 8; period: 2π; midline: y = 5 Amplitude: 4; period: 2π; midline: y = 5 Amplitude: 2; period: π; midline: y = 1

OpenStudy (misty1212):

really need the graph for this one

OpenStudy (itsmichelle29):

okay =)

OpenStudy (misty1212):

ok much better got a guess?

OpenStudy (misty1212):

only need to look at the amplitude to pick from your choices

OpenStudy (itsmichelle29):

Amplitude: 8; period: π; midline: y = 1 Amplitude: 8; period: 2π; midline: y = 5 Amplitude: 4; period: 2π; midline: y = 5 Amplitude: 2; period: π; midline: y = 1

OpenStudy (itsmichelle29):

well i thought it was -1 but thats not a chose

OpenStudy (misty1212):

from -1 to 3, that is from the lowest point to the highest, is a distance of 4 units the amplitude is half of that i.e. 2

OpenStudy (misty1212):

also you can see that the period is \(\pi\) since it goes up, and back down to -1 from 0 to \(\pi\)

OpenStudy (itsmichelle29):

Amplitude: 2; period: π; midline: y = 1 is that the answer

OpenStudy (misty1212):

yup

OpenStudy (itsmichelle29):

omg ur amazing thanks

OpenStudy (misty1212):

lol thanks\[\color\magenta\heartsuit\]